Suppose we have a market demand Q = 18 – P and a cost C(Q) 9) = 3Q?.Suppose a second firm enters the market described in question 1 (market demand is 1 still Q = 18 – P) with the same cost (cle
Answer #1

(a)

A collusion of firms act like a monopolist, equating MR and MC.

P = 18 - Q

TR = PQ = 18Q - Q2

MR = dTR/dQ = 18 - 2Q

MC = dC/dQ = Q

18 - 2Q = Q

3Q = 18

Q = 6

P = 18 - 6 = 12

(b)

TR = 12 x 6 = 72

TC = (6 x 6 / 2) = 18

Profit = TR - TC = 72 - 18 = 64

(c)

A collusion of multiple firms is called a Cartel.

(d)

A Cartel is inherently unstable, because each member of the cartel has the incentive to cheat and deviate from the cartel agreement, by either selling higher quantity or by selling at lower price.